OpenStudy (anonymous):

The weight of an object on or above the surface of Earth varies inversely as the square of the distance between the object and Earth’s center. If a person weighs 150 pounds on Earth’s surface, find the person’s weight 700 miles above the surface of the Earth. Assume that the radius of the Earth is 4000 miles. Round your answer to the nearest whole pound. HELPPPPP

OpenStudy (anonymous):

\[150 =\frac{ C }{ 4000^{2} } => x= \frac{ C }{ 4700^{2} }\]

OpenStudy (anonymous):

IS C/4700^2 THE FIANL ANSWER OR DO I HAVE TO REDUCE IT

OpenStudy (anonymous):

SOLVE LEFT SIDE FOR C USE C TO SOLVE FOR X

OpenStudy (anonymous):

IM STILL LOST COULD YOU SHOW ME

OpenStudy (anonymous):

U GOT THIS

OpenStudy (anonymous):

DO ALGEBRA ETC

OpenStudy (anonymous):

AM IM ADDING OR SUBTRACTING

OpenStudy (anonymous):

NEITHER

OpenStudy (anonymous):

SO WHAT DO I DO FROM HERE JUST LEAVE IT AS IS

OpenStudy (anonymous):

HAVE YOU LEARNED MULTIPLYING YET? IT COMES AFTER ADDING

OpenStudy (anonymous):

YEA I DID

OpenStudy (anonymous):

COOL

OpenStudy (anonymous):

SO IM DIVIDNG C/4700 THATS ALL

OpenStudy (anonymous):

WOAH HOLD ON LET'S NOT GET CRAZY HERE WHO SAID ANYTHING ABOUT DIVIDING THAT'S ADVANCED

OpenStudy (anonymous):

SOULD I BE MULTIPLYING C/4700

OpenStudy (anonymous):

SOLVE LEFT SIDE FOR C USE C TO SOLVE FOR X

OpenStudy (anonymous):

\[\huge 150 = \frac{ C }{ 4000^{2} }\]

OpenStudy (anonymous):

SOLVE THAT FOR C

OpenStudy (anonymous):

24000000

OpenStudy (anonymous):

NEEDS MORE ZEROES